Estimate the number of basketballs it would take to fill your classroom (for distance students your “classroom” can be whatever room you use to watch the lectures). State any assumptions you make. Specify how many significant figures you think are appropriate for your answer.
Answer #1

Assume the dimensions of the room

Length L = 10 ft

Width W = 10 ft

Height H = 10 ft

Volume of room V = L x W x H

= 10 ft x 10 ft x 10ft

= 1000 ft3

Dia of basket ball D = 9.55 in x (1ft/12in) = 0.79583 ft

Volume of basket ball Vb = (4/3)*3.14*(D/2)^3

= (4/3)*3.14*(0.79583/2)^3

= 0.26378 ft3

Number of basketballs = V/Vb

= 1000/0.26378

= 3791

4 significant figures are sufficient to deal with this problem because number of basketballs can be fit in a room is in thousands.